OpenStudy (anonymous):

The sum of the squares of two consecutive positive even numbers is 52. What is the smaller number? @mathstudent55 @noemy_bieber

OpenStudy (anonymous):

x^2+(x+2)^2=52 x^2+x62+4x+4=52 2x62+4x+4-52=0 2x^2+4x-48=0 2(x^2+2x-24)=0 2(x+6)(x-4)=0 x+6=0 x=-6 ans. x-4=0

OpenStudy (anonymous):

x=4 ans. Proof: -6^2+(-6+2)^2=52 36-4^2=52 36+16=52 52=52 4^2+(4+2)^2=52 16+6^2=52 16+36=52 52=52

OpenStudy (anonymous):

is it 4 or -6

OpenStudy (dan815):

hmm ur equation doesnt necessaliry say those are even numbers but ill let it pass..

OpenStudy (anonymous):

im not sure

OpenStudy (dan815):